# Coverage Axis++: Efficient Inner Point Selection for 3D Shape Skeletonization

### v5

@@ Abstract.
We introduce Coverage Axis++, a novel and efficient approach to 3D shape skeletonization. The current state-of-the-art approaches for this task often rely on the watertightness of the input (Li et al., 2015) or suffer from substantial computational costs (Dou et al., 2022), thereby limiting their practicality. To address this challenge, Coverage Axis++ proposes a heuristic algorithm to select skeletal points, offering a high-accuracy approximation of the Medial Axis Transform (MAT) while significantly mitigating computational intensity for various shape representations. We introduce a simple yet effective strategy that considers both shape coverage and uniformity to derive skeletal points. The selection procedure enforces consistency with the shape structure while favoring the dominant medial balls, which thus introduces a compact underlying shape representation in terms of MAT. As a result, Coverage Axis++ allows for skeletonization for various shape representations (e.g., water-tight meshes, triangle soups, point clouds), specification of the number of skeletal points, few hyperparameters, and highly efficient computation with improved reconstruction accuracy. Extensive experiments across a wide range of 3D shapes validate the efficiency and effectiveness of Coverage Axis++. The code will be publicly available once the paper is published.
medial axis transform, medial surface, skeletonization, shape analysis
@@
submissionid: ***ccs: Computing methodologies Shape analysisccs: Computing methodologies Point-based modelsccs: Computing methodologies Mesh geometry models
1. Introduction
@@
Skeletal representations have become a popular tool in various applications of shape analysis and geometric processing, as they efficiently capture the underlying structures of 3D shapes. Skeletal representations have been widely adopted for various downstream tasks including 3D reconstruction (Wu et al., 2015; Tang et al., 2019; Amenta et al., 2001), volume approximation (Stolpner et al., 2011; Sun et al., 2013), shape segmentation (Lin et al., 2020), shape abstraction (Dou et al., 2020), pose estimation (Shotton et al., 2011; Yang et al., 2021), and animation (Baran and Popović, 2007; Yang et al., 2018), among others.
@@
Previous efforts have been made for the computation of curve skeleton (Au et al., 2008; Ma et al., 2003; Tagliasacchi et al., 2012; Xu et al., 2019), which consists of only 1D curves. While Dey and Sun (Dey and Sun, 2006) provide a mathematical definition based on the Medial Geodesic Function, curve skeletons are generally empirically understood. Moreover, the curve representation is typically limited to tubular components and cannot be considered a generalized tool for shape analysis.
@@
Another popular skeletal representation is the Medial Axis Transform (MAT) (Blum et al., 1967). The MAT is defined by a union of maximally inscribed balls within the shape, along with their associated radius functions. A formal definition of MAT is given in Sec. 3.1. Different from the curve skeleton, the MAT has a consistent definition for arbitrary shapes. It comprises both curve-like and surface-like structures, leading to a significantly better representational ability.
@@
As of yet, the MAT computation poses challenges due to sensitivity to boundary noise and strict input geometry requirements, e.g., watertightness and manifoldness of the surface (Li et al., 2015). To tackle the problem, an effective method named Coverage Axis (Dou et al., 2022) is introduced to model skeletal point selection as a Set Cover Problem. It aims to identify the smallest sub-collection covering the entire shape. This method minimizes dilated inner balls to approximate the shape, establishing a correspondence between the skeleton and the shape, leading to a compact representation of the original shape. However, solving the Coverage Axis is time-consuming since it is in an SCP format, which is a well-known NP-hard problem. For instance, the average running time of the Coverage Axis can be $236.1$s with the standard deviation being $408.4$s; See Tab. 1.
@@
In this paper, we develop a simple yet more efficient method for shape skeletonization. We take centrality and uniformity into consideration and sidestep the expensive coverage constraint by developing a heuristic algorithm to achieve representative skeletal point selection with high efficiency. Specifically, we present a scoring scheme that assigns a score to each inner point as a quantification of its skeletal representation ability. The scoring consists of the coverage score and the uniformity score. The coverage score is determined by evaluating the number of surface samples that can be covered within the dilated ball associated with the candidate to ensure coverage. The coverage score enforces consistency with the shape structure while favoring the interior points that dominate a larger area, which corresponds to the definition of MAT, i.e., maximally inscribed spheres. The centrality of the skeleton is achieved by the coverage condition. On the other hand, uniformity is one of the key desired features of compact skeletonization without redundant structures, for which we introduce the uniformity score. The score is calculated by measuring the distance between the candidate and the nearest point in the set of selected inner points. This achieves a uniformly distributed skeletonization result by preventing the selected points from clustering together. The skeletal points are derived by running a priority queue defined by these two scores. This simple yet effective point selection manner produces a better abstraction of the overall shape as well as a shape-aware point distribution and sidesteps solving an optimization problem with high computation complexity. Notably, although the priority queue works in a greedy manner, Coverage Axis++ achieves global consistency between the skeleton and the original shape and allows for specifying different numbers of target skeletal points; See Sec. 6.1.
@@
Compared with the existing methods (Amenta et al., 2001; Li et al., 2015; Giesen et al., 2009; Foskey et al., 2003; Dey and Zhao, 2002; Sud et al., 2007; Miklos et al., 2010; Dou et al., 2022), this simple yet effective strategy allows for 1) skeletonization for various shape representations including water-tight meshes, triangle soups and point clouds; 2) specifying the number of skeletal points, which is a desirable feature for learning-based methods as a fixed number of points are typically required; 3) highly efficient computation while achieving better or more competitive reconstruction accuracy; 4) few hyperparameters; 5) randomly generated candidates inside the volume as it selects the most expressive ones from overfilled inner point candidates.
@@
We conduct extensive experiments to demonstrate the effectiveness and robustness of our method on a variety of 3D shapes. The comprehensive evaluation reveals that our method effectively captures the underlying MATs with lower reconstruction errors while achieving much higher computation efficiency.
2. Related Work
@@
Efficient shape representation is essential for various geometric modeling and analysis tasks (Amenta et al., 2001; Marschner et al., 2023; Liu et al., 2023b; Sellán et al., 2020; Xu et al., 2022, 2023; Dou et al., 2020; Xin et al., 2016; Fu et al., 2022; Sellán et al., 2021; Wen et al., 2023; Lin et al., 2020, 2023; Yang et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Yariv et al., 2023; Yu et al., 2023). Skeletonization, as a popular choice in this regard, has been extensively studied in recent decades. We refer readers to (Tagliasacchi et al., 2016) for a detailed survey covering various forms of skeletons.
Curve Skeletonization
Many efforts have been made to curve skeletonization. Traditional approaches typically rely on hand-crafted rules to leverage geometric information to compute curved skeletons (Attali and Montanvert, 1996; Ma et al., 2003; Sharf et al., 2007; Au et al., 2008; Natali et al., 2011; Tagliasacchi et al., 2012; Xu et al., 2019; Cheng et al., 2020; Bærentzen and Rotenberg, 2021). Ma et al.(Ma et al., 2003) employ radial basis functions (RBFs) for extracting the skeleton. Sharf et al. (Sharf et al., 2007) utilize a deformable model evolution that captures the volumetric shape of the object, subsequently generating an approximation for the curve skeleton. Au et al.(Au et al., 2008) compute the skeleton through mesh extraction. Tagliasacchi et al. (Tagliasacchi et al., 2012) formulate the skeletonization problem using mean curvature flow (MCF) and incorporate external energy terms to generate the curve skeleton.
Livesu et al.(Livesu et al., 2012) use the visual hull for 3D shape curve skeleton reconstruction. Mean Curvature Skeleton(Tagliasacchi et al., 2012) employs mean curvature flow to collapse the input mesh, generating a meso-skeleton with surface-like and curve-like structures. Cheng et al. (Cheng et al., 2020) propose skeletonization using the dual of shape segmentation. Learning-based methods, as proposed in (Xu et al., 2019), aim to predict curve skeletons. However, these skeletons face challenges in accurately representing arbitrary shapes, particularly those with flat components.
@@
The Medial Axis Transform (MAT), as a more general skeletal representation, encodes shapes with curve-like and mesh-like structures by representing them through maximal balls with varying radii. It can be viewed as a discrete Minkowski sum, providing a concise mathematical description of the shape.
@@
Efforts in MAT computation have been extensive (Amenta et al., 2001; Foskey et al., 2003; Dey and Zhao, 2002; Sud et al., 2007; Hu et al., 2023; Ge et al., 2023; Hu et al., 2022; Dou et al., 2022; Rebain et al., 2021; Wang et al., 2022a). Angle-based filtering methods (Amenta et al., 2001; Foskey et al., 2003; Dey and Zhao, 2002; Sud et al., 2007) simplify based on the angle formed by a point of MA and its two closest boundary points, preserving local features but compromising original topology. $\lambda$-medial axis methods (Chazal and Lieutier, 2005; Chaussard et al., 2011) use cumradius of closest points as a pruning criterion but struggle with feature preservation at different scales (Attali et al., 2009). Scale Axis Transform (Miklos et al., 2010) excels in spike pruning but is limited by high computational cost and topology disruption. Progressive medial axis (Faraj et al., 2013) proposes to perform successive edge collapse based on sphere absorption to preserve the topology as well as improve time efficiency. Besides, many strategies have been proposed for MAT simplification including Delta Medial Axis (DMA)(Marie et al., 2016), Bending Potential Ratio pruning(Shen et al., 2011), Erosion Thickness measure (Yan et al., 2016), and voxelization-based $\lambda$-pruning(Yan et al., 2018). Sun et al.(Sun et al., 2013) propose to control the approximation error through Hausdorff distance. Li et al. (Li et al., 2015) propose Q-MAT for highly accurate shape approximation, utilizing a quadratic error metric (Garland and Heckbert, 1997) during simplification. However, Q-MAT is limited to 2-manifold surfaces. Besides, Q-MAT is sensitive to the quality of MAT initialization. Recently, Wang et al. (Wang et al., 2022a) compute MAT with well-preserved geometric features via the Restricted power diagram.
Point Cloud Skeletonization
Point cloud skeletonization has gained attention due to the widespread availability of point cloud data (Andrea Tagliasacchi, 2009; Cao et al., 2010; Livesu et al., 2012; Huang et al., 2013; Rebain et al., 2019; Wu et al., 2015; Yang et al., 2020). The L1-medial skeleton (Huang et al., 2013) contracts point clouds using locally optimal projections (Lipman et al., 2007). LSMAT(Rebain et al., 2019) approximates the MAT based on the Signed Distance Function from a densely sampled oriented point set. Wu et al.(Wu et al., 2015) link surface points to the inner point on the meso-skeleton(Tagliasacchi et al., 2012), but their unstructured point representation lacks topological constraints, leading to high reconstruction errors. With the rise of deep learning, learning-based skeletonization methods for point clouds have emerged (Yang et al., 2020; Lin et al., 2021; Ge et al., 2023). Point2Skeleton (Lin et al., 2021) learns skeletal representations from generalized point clouds. However, learning-based methods may lack accuracy in computing geometric features and face challenges in generalization due to dependency on training data as we will demonstrate.
3. Preliminaries
3.1. Medial Surface
@@
Figure 2. The pipeline of Coverage Axis++. (a) Input point cloud. (b) Inner point candidates. (c) Candidate inner balls with original radius. (d) Selected inner balls with original radius. (e) Selected inner points. (f) Skeletonization Result.
3.2. Voronoi Diagram & Power Diagram
@@
A well-known MAT initialization technique is to use the VD inside a shape generated by surface samples (Amenta et al., 2001; Li et al., 2015; Dou et al., 2022).
In Power Diagram (PD) (Aurenhammer, 1987), each generator $\mathbf{x}_{i}$ is equipped with a weight $w_{i}$ to control its influence. By defining the power distance $d^{\text{pow}}(x,x_{i})$ between $x$ and the weighted generator $x_{i}$ to be $\|\mathbf{x}-\mathbf{x}_{i}\|^{2}-w_{i}$, the cell associated with $\mathbf{x}_{i}$ is defined by
@@
A generator with a larger weight is more dominant, and when weights are equal, the PD is reduced to a VD.
3.3. Coverage Axis
Coverage Axis (Dou et al., 2022) is an effective method for shape skeletonization of both meshes and point clouds, where the skeletal point selection problem is modeled as a set cover problem (SCP). Given candidate skeletal points $P=\{p_{i}\}$, the goal is to find the minimum number of dilated balls that cover all the sampled points $S=\{s_{j}\}$ on the surface. A coverage matrix $\mathbf{D}\in\{0,1\}^{m\times n}$ is used, where $m$ and $n$ are total numbers of sampled surface points and candidate skeletal points, respectively. Each entry $d_{ji}\in\{0,1\}$ of $\mathbf{D}$ indicates if the surface point $s_{j}$ is covered by the dilated ball $(p_{i},r_{i}^{\prime})$:
(1) $d_{ji}=\left\{\begin{aligned} 1,&\ {\color[rgb]{0,0,0}\text{if}}\ \left\|p_{i}-s_{j}\right\|_{2}\leq r_{i}^{\prime},j=1,...,m,\ i=1,...,n\\ 0,&\ {\color[rgb]{0,0,0}\text{if}}\ \left\|p_{i}-s_{j}\right\|_{2}>r_{i}^{\prime},j=1,...,m,\ i=1,...,n.\\ \end{aligned}\right.$
Let $\mathbf{v}\in\{0,1\}^{n\times 1}$ be a decision vector, where the $i$-th element $v_{i}\in\{0,1\}$ indicates if the candidate skeletal point $p_{i}$ is selected. Coverage Axis selects points by solving the following constrained 0-1 integer programming problem:
(2) $\begin{split}&\min\left\|\mathbf{v}\right\|_{2}\\ &s.t.\ \begin{array}[]{c}\mathbf{Dv}\geq\mathds{1},\\ \end{array}\end{split}$
where $\mathds{1}$ is the vector of all ones and $\geq$ is applied element-wise. The neat formulation achieves high-level abstraction and robustness to noise. However, SCP is a well-known NP-Hard problem (Hartmanis, 1982) and its running time complexity is $O(|P|^{2}2^{|S|})$ where $|P|$ and $|S|$ are the number of candidate inner points and surface points, respectively. As a result, the Coverage Axis suffers from a high computational cost; Refer to Tab. 1 and Appendix. A.3.
4. Method

### v7

@@ Abstract
We introduce Coverage Axis++, a novel and efficient approach to 3D shape skeletonization. The current state-of-the-art approaches for this task often rely on the watertightness of the input [LWS15, PWG19, PWG19] or suffer from substantial computational costs [DLX22, CD23], thereby limiting their practicality. To address this challenge, Coverage Axis++ proposes a heuristic algorithm to select skeletal points, offering a high-accuracy approximation of the Medial Axis Transform (MAT) while significantly mitigating computational intensity for various shape representations. We introduce a simple yet effective strategy that considers shape coverage, uniformity, and centrality to derive skeletal points. The selection procedure enforces consistency with the shape structure while favoring the dominant medial balls, which thus introduces a compact underlying shape representation in terms of MAT. As a result, Coverage Axis++ allows for skeletonization for various shape representations (e.g., water-tight meshes, triangle soups, point clouds), specification of the number of skeletal points, few hyperparameters, and highly efficient computation with improved reconstruction accuracy. Extensive experiments across a wide range of 3D shapes validate the efficiency and effectiveness of Coverage Axis++. Our codes are available at https://github.com/Frank-ZY-Dou/Coverage_Axis.
{CCSXML}
@@
<ccs2012> <concept> <concept_id>10010147.10010371.10010396.10010402</concept_id> <concept_desc>Computing methodologies Shape analysis</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396.10010400</concept_id> <concept_desc>Computing methodologies Point-based models</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396.10010398</concept_id> <concept_desc>Computing methodologies Mesh geometry models</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
volume: 43issue: 5
1 Introduction
@@
Skeletal representations have become a popular tool in various applications of shape analysis and geometric processing, as they efficiently capture the underlying structures of 3D shapes. They have been widely adopted for various tasks including 3D reconstruction [WHG15, THP19, ACK01], volume approximation [SKS11, SCYW13], shape segmentation [LLL20], shape abstraction [DXX20], pose estimation [SFC11, YRL21, DWL23, LXC21], and animation [BP07, YYG18], among others. With the development of deep learning, skeletal representations also facilitate learning-based approaches in various tasks [HWQ19, RLS21, THT21, HCY22, PGT24, WYT23, XZX24, WTJL24, GWHM24].
@@
Previous efforts have been made for the computation of curve skeleton [ATC08, MWO03, TAOZ12, XZKS19], which consists of only 1D curves. Dey and Sun [DS06] provide a mathematical formulation based on the Medial Geodesic Function, yet curve skeletons are predominantly comprehended empirically.
@@
Medial Axis Transform (MAT) [B67], as a complete shape descriptor, is another popular skeletal representation. The MAT is defined by a union of maximally inscribed balls within the shape, accompanied by their respective radius functions. A formal definition of MAT can be found in Sec. 3.1. In contrast to curve skeletons, the MAT encompasses both curve-like and surface-like structures, offering a consistent definition for arbitrary shapes and superior representational capability.
@@
The MAT computation poses challenges due to sensitivity to boundary noise [LWS15] and strict input geometry requirements, e.g., watertightness and manifoldness of the surface [LWS15, PWG19, WWWG22]. To tackle the problem, Coverage Axis [DLX22] is introduced to model skeletal point selection as a Set Cover Problem (SCP), which does not rely on surface connectivity and is able to handle both mesh or point cloud inputs. It aims to identify the smallest sub-collection covering the entire shape. This method minimizes dilated inner balls to approximate the shape, leading to a compact representation of the original shape. However, solving the SCP problem, which is known to be an NP-hard problem, can be time-consuming in some cases. For instance, the average running time of the Coverage Axis can be $236.1$s with the standard deviation being $408.4$s; See Tab. 1. Meanwhile, the recent work Neural Skeleton [CD23] proposes to first learn a Signed Distance Field from the given shape and then use the Coverage Axis to get the final compact skeleton of the given mesh or point cloud. However, it still suffers from low runtime performance (59.27s compared to our 1.63s) and relatively low reconstruction accuracy as shown in Tab. 3.
@@
In this paper, we develop a simple yet more efficient skeletonization method for shapes represented as meshes or point clouds. We take shape representation capability, centrality, and uniformity into consideration and develop a heuristic algorithm to achieve representative skeletal point selection with high efficiency. Specifically, we present a scoring scheme that assigns a score to each inner point as a quantification of its skeletal representation ability. The scoring consists of the coverage, uniformity and centrality scores. The coverage score is determined by evaluating the number of surface samples that can be covered within the dilated ball associated with the candidate to ensure coverage inspired by  [DLX22]. The coverage score ensures consistency with the shape structure by favoring interior points that dominate a larger area, aligning with the definition of the Medial Axis Transform (MAT) as the set of centers of maximally inscribed spheres. On the other hand, uniformity is one of the key desired features of compact skeletonization to suppress redundant structures, for which we introduce the uniformity score. The score is calculated by measuring the distance between the candidate and the nearest point in the set of selected inner points. This achieves a uniformly distributed skeletonization result by preventing the selected points from clustering together. Moreover, the centrality score is employed to encourage those central points to receive higher scores to promote centrality further. The skeletal points are derived by running a priority queue defined by an integration of these scores. We show that this simple yet effective point selection manner produces a better abstraction of the overall shape as well as a shape-aware point distribution without the need to solve an optimization problem with high computation complexity. Notably, although the priority queue works in a greedy manner, Coverage Axis++ achieves global consistency between the skeleton and the original shape and allows for specifying different numbers of target skeletal points; See Sec. 6.1.
@@
Compared with the existing methods [ACK01, LWS15, GMPW09, FLM03, DZ02, SFM07, MGP10, DLX22], Coverage Axis++ surpasses all existing methods in terms of both MAT approximation accuracy and computation efficiency while being able to handle various inputs, including mesh, point cloud or polygon soup. To summarize, our method allows for 1) skeletonization for various shape representations including water-tight meshes, triangle soups and point clouds; 2) specifying the number of skeletal points, which is a desirable feature for learning-based methods as a fixed number of points are typically required [HWQ19, THT21, HCY22, PGT24]; 3) highly efficient computation while achieving better or more competitive reconstruction accuracy; 4) few hyperparameters, i.e., only dilation factor is required; 5) randomly generated candidates inside the volume as Coverage Axis++ selects the most expressive ones from overfilled inner point candidates.
@@
We conduct extensive experiments to demonstrate the effectiveness and robustness of our method on a variety of 3D shapes. The comprehensive evaluation reveals that our method effectively captures the underlying MATs with lower shape approximation errors while achieving much higher computation efficiency.
2 Related Work
@@
Efficient shape representation is essential for various applications of geometric modeling, shape analysis and shape generation [MSLJ23, LGC23, SKSJ20, XWD22, XDW23, DXX20, FXX22, WYT23, LLL20, LYZ23, YLL23, ZLY23, YPN23, YDL23, Ale21, PGT24]. Skeletonization, as a popular choice in this regard, has been extensively studied in recent decades. We refer readers to [TDS16] for a detailed survey covering various forms of skeletons. In the following, we mainly review the related works in Medial Axis Transform (MAT) computation for mesh and point cloud inputs.
@@
Mesh Inputs. Efforts in MAT computation given the watertight mesh inputs have been extensive [ACK01, GWM23, FLM03, DZ02, SFM07, HWY23, GYY23, HCY22, DLX22, RLS21, WWWG22, LLL24, LFWN20]. Specifically, $\lambda$-medial axis [CL05] define weak feature size as a pruning criterion but struggle with feature preservation at different scales [ABE09]. Scale Axis Transform [MGP10] excels in spike pruning but is limited by high computational cost and topology disruption. Progressive medial axis [FTB13] proposes to perform topology-preserved edge collapse based on sphere absorption. Many studies have been conducted for MAT simplification using different metrics [MLIM16, SBH11, YSC16, YLJ18, SCYW13, LWS15, PWG19]. Li et al. [LWS15] propose Q-MAT for highly accurate shape approximation, utilizing a quadratic error metric [GH97] during simplification. Despite relatively high approximation accuracy, Q-MAT relies on watertight surfaces. It is sensitive to the quality of MAT initialization, as revealed by  [DLX22]. Later, Q-MAT+ [PWG19], as a variant of Q-MAT, further introduces more mesh information like shape diameter function to improve the approximation accuracy and robustness of the original Q-MAT method, especially for those sharp features. Recently, MATFP [WWWG22] is developed to compute MAT with well-preserved geometric features via the Restricted Power Diagram for CAD mesh models. Wang et al. [WHS24] have also explored addressing the topology preservation issue by employing Restricted Power Diagrams.
Point Cloud Inputs. Point cloud skeletonization is still a fascinating yet challenging problem [AT09, CTO10, LGS12, LZG23, HWCO13, RAV19, WHG15, YYW20, DLX22] due to the absence of manifold information regarding the underlying surface of the shape. Specifically, L1-medial skeleton [HWCO13] employs Locally Optimal Projection [LCOLTE07] to contract point clouds to form the skeleton. LSMAT[RAV19] approximates the MAT based on the Signed Distance Function (SDF) from a densely sampled oriented point set. Wu et al.[WHG15] considers linking surface points to the skeletal points on the meso-skeleton[TAOZ12]. Yet, this representation only yields unstructured point sets, which lacks topological constraints describing the skeleton connectivity and leads to low approximation accuracy. Coverage Axis [DLX22] models skeletal point selection as a Set Cover Problem (SCP), which does not rely on surface connectivity, thus allowing for point cloud inputs. However, it suffers from high computational costs as SCP is NP-hard, especially when facing shapes with planar structures (See detailed analysis in Appendix D).
With the advent of deep learning, learning-based approaches have been explored for point clouds skeletonization [YYW20, LLL21, GYY23]. For example, Point2Skeleton [LLL21] learns to predict a set of skeletal points by learning a geometric transformation and then analyzes the connectivity of the skeletal points to form mesh structures from point clouds. However, it suffers from limited generalization due to dependency on training data and lacks accuracy for geometric features. Recently, Neural Skeleton [CD23] leverages Implicit Neural Representations (INR), i.e., SDF, together with coverage axis [DLX22] for skeleton computation, yet it still faces challenges with long running time and dependence on consistently oriented normals for input shapes.
3 Preliminaries
3.1 Medial Axis Transform
@@
Figure 1: The pipeline of Coverage Axis++. (a) Input point cloud. (b) Inner point candidates. (c) Candidate inner balls with original radius. (d) Selected inner balls with original radius. (e) Selected inner points. (f) Skeletonization result.
3.2 Voronoi Diagram & Power Diagram
@@
A well-known MAT initialization technique is to use the VD inside a shape generated by surface samples [ACK01, LWS15, DLX22].
In Power Diagram (PD) [Aur87], each generator $\mathbf{x}_{i}$ is equipped with a weight $w_{i}$ to control its influence. By defining the power distance $d^{\text{pow}}(x,x_{i})$ between $x$ and the weighted generator $x_{i}$ to be $\|\mathbf{x}-\mathbf{x}_{i}\|^{2}-w_{i}$, the cell associated with $\mathbf{x}_{i}$ is defined by
@@
Therefore, a generator with a larger weight is more dominant.
4 Method